# Strain Problems got you in a Twist? Try StrainRelief: A Quantum-Accurate Tool for Ligand Strain Calculations

### v1

@@ 1 Introduction
Structure-based drug design has been an incredibly effective strategy for drug discovery since structural biology techniques like x-ray crystallography entered the field in the 1980s.1 In this paradigm, new drugs are designed by chemists to fit into an experimentally determined, 3D protein pocket. This “lock and key” style approach intimately relies on accurately predicting and evaluating the pose of any proposed ligand molecule in its pocket. One of the most important metrics in evaluating the quality of a ligand pose is the ligand strain energy.2 Strain energy is defined as the energy difference between the bound and unbound conformations of a ligand (Eqn. 1), and this metric is valuable because the vast majority of ligand conformations observed in experimentally measured drug-protein complexes have low strain.2 The logic of this is simple. A ligand will only bind to a protein if it is energetically favourable for it to do so. If the conformation required for a ligand to fit into a pocket is too strained, it will not bind. This concept is not just true in the abstract; ligand strain energy has been used as a metric to drive success on real drug discovery projects.3, 4 An accurate ligand strain energy filter can have great value for structure-based drug discovery. In this paper, we present StrainRelief, a tool that measures ligand strain energy near the accuracy of ab initio quantum chemistry.
@@
All of this is combined into the StrainRelief tool, which we make publicly available here.
@@ 2.1 Accurate Energy Evaluation
In this study, a selection of NNPs, including MACE20, ANI21, 22, and NequIP23, along with a commonly used empirical potential, MMFF94s,24 were evaluated on the local and global minimum conformers of our own curated neutral subset of the LigBoundConf dataset2, LigBoundConf 2.0 (see Methods - LigBoundConf 2.0) (Figure 2). All NNPs were trained on neutral molecules from SPICE 225 - a dataset curated to encapsulate protein-ligand interactions with maximally diverse conformers. Conformers are evaluated with DFT at the $\omega$B97M-D3(BJ)26, 27/def2-TZVPPD28, 29 level of theory. This level of theory is chosen for its balance of speed and accuracy on quantum chemistry benchmarks.30 To unlock significantly higher accuracy a double hybrid functional or quadruple zeta basis set would be needed, both of which are prohibitively expensive for a dataset of this size. Among the models tested, MACE demonstrated superior performance, exhibiting a significantly lower mean absolute error (MAE). MACE is roughly equally accurate in evaluating bound and global minimum conformers with MAEs of SPE predictions of 1.85 and 1.72 kcal/mol respectively. Crucially, a Spearman’s rank correlation of 0.92 indicates that MACE effectively identifies the strain of ligands relative to one another - an important quality for a filter. Despite its overall strong performance, MACE tends to perform poorly in certain cases, with predictions deviating by as much as 10 or 20 kcal/mol. Zwitterionic ligands are particularly problematic, exhibiting an average MAE that is 4.7 kcal/mol higher than other ligands. These zwitterionic ligands account for 62% of cases with errors above 10 kcal/mol and 57% of cases with errors above 20 kcal/mol. This can be attributed to the fact that charged molecules in the gas phase tend to form unphysical shapes, and all charged molecules were excluded from the SPICE 2 training data.
@@
MACE generalises well to LigBoundConf 2.0, with only 16 ligands being common between this and the SPICE 225 training set. Excluding common ligands, a further 2 ligands are considered to be highly structurally similar to any molecules in the SPICE 225 training set, determined by a minimum Tanimoto difference $\leqslant 0.15$ (Figure S1).31 MACE’s exceptional performance is consistent with findings from recent literature on this NNP architecture17, 32, 33, 34, 35. MACE achieves a MAE of 1.32 kcal/mol (0.068 kcal/mol per atom) on the SPICE 2 PubChem test set when predicting absolute energies. This result aligns with models trained with identical hyperparameters in MACE-OFF2317, which report a MAE of 0.063 kcal/mol per atom on SPICE 1 PubChem. Furthermore, MACE aligns closely with the MACE-OFF23 (S) model on the TorsionNet dataset and shows considerable improvements over other models presented in this study (Figure S2). Due to its excellent performance, as show in Figure 2, MACE has been selected as the standard model for StrainRelief and is utilised throughout this paper, unless specified otherwise.
@@
In addition to evaluating given conformers, StrainRelief seeks to accurately approximate minima. We ran the StrainRelief protocol with our trained MACE model on the LigBoundConf 2.0 dataset to obtain both local and global minima (Figure 3). We then evaluated the energy difference of those minima with DFT to assess the accuracy of our model+protocol. We did the same experiment with the MMFF94s model as a baseline. All minimisations were run with the Broyden–Fletcher–Goldfarb–Shanno36 (BFGS) algorithm as implemented in Atomic Simulation Environment37 (ASE) using their MACE calculator.
@@ 2.3 Enrichment of Active Samples
The MM strain15 tool is a chemalot command-line workflow (sdfMMConfAnalysis) and is the most similar of the two literature methods to StrainRelief. It calculates a local minimum and performs a global minimum search by enumerating conformers. Two key differences are that both an implicit solvent model and an empirical force field (MMFF94s) are used.
@@ 3 Methods
To obtain the “bound” structures we sought to relieve any artificial high-energy artifacts from the deposited PDB structures. We did this by running a local minimisation with the semi-empirical quantum chemistry method GFN2-xTB using the “crude” convergence criteria. This ensured that artifacts were removed without significantly altering the bound pose of the ligand. The resulting minimised structures have an average RMS displacement of 0.20 Å (95th percentile = 0.47 Å) relative to the “raw” structures. We then performed SPE DFT calculations at the $\omega$B97M-D3(BJ)26, 27/def2-TZVPPD28, 29 level of theory.
To obtain the “global minimum” structures, we used the CREST enumeration method with 50 conformers, which uses the same GFN2-xTB energy function to search for low energy gas phase structures. We took the lowest energy structure as the “global minimum” conformation and performed a SPE DFT calculation at the same $\omega$B97M-D3(BJ)26, 27/def2-TZVPPD28, 29 level of theory.
Note that this protocol is only strictly valid for neutral molecules. Charged molecules are liable to undergo electrostatic collapse in the gas phase. Moreover, including a solvation model in strain energy is troublesome; bound ligands are typically at least partially buried in protein pockets, so including solvation in a reference calculation for the bound structure is unlikely to be a useful model.
The LigBoundConf 2.0 neutral dataset ended up with 3132 ligands. The strain energy of these ligand’s difference in DFT energies between the “bound” and “global minimum” conformations is plotted in the first histogram of Figure 4. This clearly shows that drug-like ligands are likely to bind with low strain energy, in agreement with the original LigBoundConf dataset. The 95th percentile of the distribution lies at 16.1 kcal/mol, which we adopted as our strained vs. unstrained classification cutoff. We use this cutoff as the classification mechanism for the subsequent experiments in this work.
@@ 3.2 StrainRelief Protocol
The protocol used in StrainRelief is designed to be simple, fast and model agnostic - all that is needed to apply a new force field is to write an ASE37 calculator wrapper. The protocol consists of 5 steps:
@@
Minimisation: All minimisations are done using an adapted version of the BFGS36 algorithm as implemented in ASE,37 with default values for $max\ step$ (0.2 Å) and $alpha$ (70.0). Minimisation is killed if the number of steps exceeds 250 or if during any step the $max(force)>250\ eV/$Å. ASE calculator wrappers were written for both MACE and RDKit’s MMFF94 force fields with the intention that other calculators could be easily written in a similar fashion. i) Local Minimisation: The given docked pose is minimised with a loose convergence criteria ($max(forces)<0.50\ eV/$Å). Determining a local minimum with your chosen force field is necessary to “clean-up” ligand structures. Docked poses can often contain artificially high energy artifacts, such as slight deviations in ideal bond lengths or angles. ii) Global Minimisation: All generated conformers (and the docked pose) are minimised with the default BFGS convergence criteria of $max(forces)<0.05\ eV/$Å.
@@
3.3 The SPICE Dataset25
All NNPs in this paper were trained on the SPICE 2.0.1 dataset. SPICE is a quantum chemistry dataset specifically designed for training potentials relevant to simulating drug-like small molecules and their interactions with proteins. SPICE conformers are generated by running molecular dynamics simulations and sampling maximally different conformations. These conformers are then evaluated at the $\omega$B97M-D3(BJ)26, 27/def2-TZVPPD28, 29 level of theory. SPICE contains a number of subsets, with PubChem being the largest, accounting for 70% of all conformers. SPICE is a useful training set for a ligand strain model because it samples many and varied high- and low-energy conformers across the drug-like small molecule space. For our training, we excluded the ion pairs subset, removed all metal-containing compounds (leaving only H, C, N, O, F, P, S, Cl, Br, and I) and removed all molecules with a non-zero formal charge. This leaves 1,834,342 conformers. We trained on an 80-10-10 train-test-validation split, stratified by subset.
@@ 3.4.1 MACE20
The MACE model was trained using the mace repository38 over 1500 epochs. Following the recommendations in the MACE-OFF23 paper, training was initially weighted towards forces with a learning rate of 0.01 and then switched to weighting towards energies for the final 20% of training epochs17 with a decreased learning rate of 0.00025. The Adam optimiser with Amsgrad was used and the exponential moving average of the weights was taken in each training step.
3.4.2 ANI21, 22
ANI is an equivariant deep neural network that uses a highly modified version of the Behler and Parrinello symmetry functions to build single-atom atomic environment vectors as molecular representations. The ANI-2x architecture was trained using the open-source PhysicsML repository39.
3.4.3 NequIP23
NequIP is an early equivariant graph neural network iteratomic potential. As recommended in the original paper when training small-molecule systems, we used 5 iteration blocks, a learning rate of 0.01 and a batch size of 5. Training for NequIP was conducted using the open-source PhysicsML repository39.
@@ 4.3 Conflicts of Interest
Figure 7: The minimum Tanimoto difference between each ligand in LigBoundConf 2.0 and all unique molecules in the SPICE 225 training data. This excludes the 16 ligands with a minimum distance of 0 (i.e. they belong to both datasets). 99% of ligands have a minimum Tanimoto difference ¿ 0.25 and the average difference is 0.47.
Figure 8: TorsionNet40 error metrics evaluated by different force fields (NequIP, MMFF94s, ANI, MACE, TorsionNet). Model predictions were compared against DFT calculations at the SPICE level of theory17. Barrier height is the predicted energy between maximum and minimum energy torsions. Relative energy represents the difference in energy between the minimum energy conformer and all other conformers of the same molecules. NequIP, ANI and MACE were trained on SPICE 225 The TorsionNet result is taken from the paper where it was trained on TorsionNet.
Figure 9: Pairplot of 4 different tools and their predictions on LigBoundConf 2.0. Diagonal plots show the distribution of each tool’s predictions (as in Figure 4). Non-diagonal plots show the correlation (or lack there of) between tool predictions. Protocols largely cannot agree on rankings of ligand strains, as represented by uncorrelated plots and low Spearman’s rank correlation scores. StrainRelief and xTB+CREST has the best agreement by a significant margin.
Choice of N for Conformer Generation: In this paper, we choose to request 20 conformers for each ligand as it provides the optimal balance between speed and accuracy for high-throughput workflows. Here, we present results when 100 conformers are instead requested (StrainRelief-N100) to demonstrate the diminishing returns of this approach.
Requesting 100 conformers results in an average of 65.5 conformers being returned. Since computation time scales linearly with the number of conformers generated, this corresponds to a 4.2-fold increase in the average time per molecule, leading to an average processing time of 184 seconds per molecule.
This significant increase in computational cost yields only modest gains. 42.0% of global minima identified by StrainRelief-N100 are lower in energy than the global minima identified by XTB+CREST. This represents an improvement of 5.5%. Crucially it provides only marginal gains when used for the filter (Table 2).
Table 2: StrainRelief v.s. StrainRelief-N100 accuracy when classifying LigBoundConf 2.0 molecules as strain or unstrained. This is an extension of Table 1.
Protocol Spearman’s R Precision Recall F1 Score
StrainRelief 0.770 0.582 0.469 0.520
StrainRelief-N100 0.794 0.579 0.620 0.599

### v2

@@ 1 Introduction
Structure-based drug design has been an incredibly effective strategy for drug discovery since structural biology techniques like x-ray crystallography entered the field in the 1980s.1 In this paradigm, new drugs are designed by chemists to fit into an experimentally determined, 3D protein pocket. This “lock and key” style approach intimately relies on accurately predicting and evaluating the pose of any proposed ligand molecule in its pocket. One of the most important metrics in evaluating the quality of a ligand pose is the ligand strain energy.2 Strain energy is defined as the energy difference between the bound and unbound conformations of a ligand (Eqn. 1), and this metric is valuable because the vast majority of ligand conformations observed in experimentally measured drug-protein complexes have low strain.2 The logic of this is simple. A ligand will only bind to a protein if it is energetically favourable for it to do so. If the conformation required for a ligand to fit into a pocket is too strained, it will not bind. This concept is not just true in the abstract; ligand strain energy has been used as a metric to drive success on real drug discovery projects.3, 4 An accurate ligand strain energy filter can have great value for structure-based drug discovery. Here we define ligand strain as an enthalpy, not a free energy, as entropic contributions are not considered. In this paper, we present StrainRelief, a tool that measures ligand strain energy near the accuracy of ab initio quantum chemistry.
@@
All of this is combined into the StrainRelief tool, which we make publicly available at github.com/prescient-design/StrainRelief.
@@ 2.1 Accurate Energy Evaluation
In this study, a selection of NNPs, including MACE20, ANI21, 16, and NequIP22, along with a commonly used empirical potential, MMFF94s,23 were evaluated on the local and global minimum conformers of our own curated neutral subset of the LigBoundConf dataset2, LigBoundConf 2.0 (see Methods - LigBoundConf 2.0) (Figure 3). All NNPs were trained on neutral molecules from SPICE 224 - a dataset curated to encapsulate protein-ligand interactions with maximally diverse conformers. Conformers are evaluated with DFT at the $\omega$B97M-D3(BJ)25, 26/def2-TZVPPD27, 28 level of theory. This level of theory is chosen for its balance of speed and accuracy on quantum chemistry benchmarks.29 To unlock significantly higher accuracy a double hybrid functional or quadruple zeta basis set would be needed, both of which are prohibitively expensive for a dataset of this size. Among the models tested, MACE demonstrated superior performance, exhibiting a significantly lower mean absolute error (MAE). MACE is roughly equally accurate in evaluating bound and global minimum conformers with MAEs of SPE predictions of 1.85 and 1.72 kcal/mol respectively. Crucially, a Spearman’s rank correlation of 0.92 indicates that MACE effectively identifies the strain of ligands relative to one another - an important quality for a filter. Despite its overall strong performance, MACE tends to perform poorly in certain cases, with predictions deviating by as much as 10 or 20 kcal/mol. Zwitterionic ligands are particularly problematic, exhibiting an average MAE that is 4.7 kcal/mol higher than other ligands. These zwitterionic ligands account for 62% of cases with errors above 10 kcal/mol and 57% of cases with errors above 20 kcal/mol. This can be attributed to the fact that charged molecules in the gas phase tend to form unphysical shapes, and all charged molecules were excluded from the SPICE 2 training data.
@@
MACE generalises well to LigBoundConf 2.0, with only 16 ligands being common between this and the SPICE 224 training set. Excluding common ligands, a further 2 ligands are considered to be highly structurally similar to any molecules in the SPICE 224 training set, determined by a minimum Tanimoto difference $\leqslant 0.15$ (Figure S1).30 MACE’s exceptional performance is consistent with findings from recent literature on this NNP architecture17, 31, 32, 33, 34. MACE achieves a MAE of 1.32 kcal/mol (0.068 kcal/mol per atom) on the SPICE 2 PubChem test set when predicting absolute energies. This result aligns with models trained with identical hyperparameters in MACE-OFF2317, which report a MAE of 0.063 kcal/mol per atom on SPICE 1 PubChem. Furthermore, MACE aligns closely with the MACE-OFF23 (S) model on the TorsionNet dataset and shows considerable improvements over other models presented in this study (Figure S2). Due to its excellent performance, as show in Figure 3, MACE has been selected as the standard model for StrainRelief and is utilised throughout this paper, unless specified otherwise.
@@
In addition to evaluating given conformers, StrainRelief seeks to accurately approximate minima. We ran the StrainRelief protocol with our trained MACE model on the LigBoundConf 2.0 dataset to obtain both local and global minima (Figure 4). The resulting local minimum structures have an average RMS displacement of 0.23 Å (95th percentile = 0.54 Å) relative to the “raw” structures. We then evaluated the energy difference of those minima with DFT to assess the accuracy of our model+protocol. We did the same experiment with the MMFF94s model as a baseline. All minimisations were run with the Broyden–Fletcher–Goldfarb–Shanno35 (BFGS) algorithm as implemented in Atomic Simulation Environment36 (ASE) using their MACE calculator.
@@ 2.3 Enrichment of Active Samples
The MM strain15 tool is a chemalot command-line workflow (sdfMMConfAnalysis) and is the most similar of the two literature methods to StrainRelief. It calculates a local minimum and performs a global minimum search by enumerating conformers. Two key differences are that both an implicit solvent model and an empirical force field (MMFF94s) are used. Figure 5 demonstrates that, for 749 of LigBoundConf 2.0’s 2726 ligands, the search for the global minimum failed to find anything lower energy than the local minimum and so predicts a strain of 0 kcal/mol.
@@ 3 Methods
To obtain the “bound” structures we sought to relieve any artificial high-energy artifacts from the deposited PDB structures. We did this by running a local minimisation with the semi-empirical quantum chemistry method GFN2-xTB using the “crude” convergence criteria. This ensured that artifacts were removed without significantly altering the bound pose of the ligand. The resulting minimised structures have an average RMS displacement of 0.20 Å (95th percentile = 0.47 Å) relative to the “raw” structures. We then performed SPE DFT calculations at the $\omega$B97M-D3(BJ)25, 26/def2-TZVPPD27, 28 level of theory.
To obtain the “global minimum” structures, we used the CREST enumeration method with 50 conformers, which uses the same GFN2-xTB energy function to search for low energy gas phase structures. We took the lowest energy structure as the “global minimum” conformation and performed a SPE DFT calculation at the same $\omega$B97M-D3(BJ)25, 26/def2-TZVPPD27, 28 level of theory.
Note that this protocol is only strictly valid for neutral molecules. Charged molecules are liable to undergo electrostatic collapse in the gas phase (Figure S4). Moreover, including a solvation model in strain energy is troublesome; bound ligands are typically at least partially buried in protein pockets, so including solvation in a reference calculation for the bound structure is unlikely to be a useful model.
The LigBoundConf 2.0 neutral dataset ended up with 2726 ligands. The strain energy of these ligand’s difference in DFT energies between the “bound” and “global minimum” conformations is plotted in the first histogram of Figure 5. This clearly shows that drug-like ligands are likely to bind with low strain energy, in agreement with the original LigBoundConf dataset. The 95th percentile of the distribution lies at 16.1 kcal/mol, which we adopted as our strained vs. unstrained classification cutoff. We use this cutoff as the classification mechanism for the subsequent experiments in this work.
@@ 3.2 StrainRelief Protocol
The protocol used in StrainRelief is designed to be simple, fast and model agnostic - all that is needed to apply a new force field is to write an ASE36 calculator wrapper. The protocol consists of 5 steps:
@@
Minimisation: All minimisations are done using an adapted version of the BFGS35 algorithm as implemented in ASE,36 with default values for $max\ step$ (0.2 Å) and $alpha$ (70.0). Minimisation is killed if the number of steps exceeds 250 or if during any step the $max(force)>250\ eV/$Å. ASE calculator wrappers were written for both MACE and RDKit’s MMFF94 force fields with the intention that other calculators could be easily written in a similar fashion. i) Local Minimisation: The given docked pose is minimised with a loose convergence criteria ($max(forces)<0.50\ eV/$Å). Determining a local minimum with your chosen force field is necessary to “clean-up” ligand structures. Docked poses can often contain artificially high energy artifacts, such as slight deviations in ideal bond lengths or angles. ii) Global Minimisation: All generated conformers (and the docked pose) are minimised with the default BFGS convergence criteria of $max(forces)<0.05\ eV/$Å.
@@
3.3 The SPICE Dataset24
All NNPs in this paper were trained on the SPICE 2.0.1 dataset. SPICE is a quantum chemistry dataset specifically designed for training potentials relevant to simulating drug-like small molecules and their interactions with proteins. SPICE conformers are generated by running molecular dynamics simulations and sampling maximally different conformations. These conformers are then evaluated at the $\omega$B97M-D3(BJ)25, 26/def2-TZVPPD27, 28 level of theory. SPICE contains a number of subsets, with PubChem being the largest, accounting for 70% of all conformers. SPICE is a useful training set for a ligand strain model because it samples many and varied high- and low-energy conformers across the drug-like small molecule space. For our training, we excluded the ion pairs subset, removed all metal-containing compounds (leaving only H, C, N, O, F, P, S, Cl, Br, and I) and removed all molecules with a non-zero formal charge. This leaves 1,834,342 conformers. We trained on an 80-10-10 train-test-validation split, stratified by subset.
@@ 3.4.1 MACE20
Models were trained using PyTorch and the training scripts in the mace repository37 with identical hyperparameters to those reported for MACE-OFF23(S)17. The model was trained for 190 epochs, with a loss function combining a linear combination of energy and force predictions. Following the recommendations in the MACE-OFF23 paper, training was initially weighted towards forces with a learning rate of 0.01. It is then switched to weighting towards energies for the final 35 training epochs17 and stochastic weight averaging (SWA) was used with a decreased learning rate of 0.00025. The Adam optimiser with Amsgrad was used and the exponential moving average of the weights was taken in each training step. The exact hyperparameters used are available in the StrainRelief GitHub repository (examples/mace_training_script.sh).
3.4.2 ANI21, 16
ANI is an equivariant deep neural network that uses a highly modified version of the Behler and Parrinello symmetry functions to build single-atom atomic environment vectors as molecular representations. The ANI-2x architecture was trained using the open-source PhysicsML repository38.
3.4.3 NequIP22
NequIP is an early equivariant graph neural network iteratomic potential. As recommended in the original paper when training small-molecule systems, we used 5 iteration blocks, a learning rate of 0.01 and a batch size of 5. Training for NequIP was conducted using the open-source PhysicsML repository38.
@@ 4.3 Conflicts of Interest
Additional information detailing the choice of N for conformer generation, the distribution of Tanimoto differences between training and tests sets, additional test set metrics, and examples of charged structures undergoing electrostatic collapse (PDF).